# STG-MTL: Scalable Task Grouping for Multi-Task Learning Using Data Maps

### v1

@@ Abstract
Multi-Task Learning (MTL) is a powerful technique that has gained popularity due to its performance improvement over traditional Single-Task Learning (STL). However, MTL is often challenging because there is an exponential number of possible task groupings, which can make it difficult to choose the best one, and some groupings might produce performance degradation due to negative interference between tasks. Furthermore, existing solutions are severely suffering from scalability issues, limiting any practical application. In our paper, we propose a new data-driven method that addresses these challenges and provides a scalable and modular solution for classification task grouping based on hand-crafted features, specifically Data Maps, which capture the training behavior for each classification task during the MTL training. We experiment with the method demonstrating its effectiveness, even on an unprecedented number of tasks (up to 100).
@@ 1 Introduction
Multi-Task Learning (MTL) has emerged as a powerful technique in deep learning (Zhang & Yang, 2022; Crawshaw, 2020) that allows for joint training of multiple related tasks, leading to improved model performance compared to traditional Single-Task Learning (STL). By leveraging shared representations and knowledge across tasks, MTL enhances generalization and mitigates overfitting. Furthermore, MTL promotes faster learning of related tasks and alleviates the computational requirements of deep learning, making it particularly valuable in scenarios with limited task-specific data. That is why MTL has gained significant attention in various domains, including computer vision (Fan et al., 2017; Misra et al., 2016; Standley et al., 2020), natural language processing (Zhang et al., 2022b; Peng et al., 2020; Jin et al., 2020; Bickel et al., 2008), speech recognition (Huang et al., 2022b; Zhang et al., 2019), and healthcare (Peng et al., 2020; Bao et al., 2022; Islam et al., 2021), and has shown promising results in improving accuracy, robustness, and efficiency. However, effectively harnessing the potential of MTL poses several challenges, including the identification of optimal task groupings (Song et al., 2022; Fifty et al., 2021; Standley et al., 2020) and the management of negative interference between tasks (Sener & Koltun, 2018; Wu et al., 2020; Maninis et al., 2019).
@@
The task grouping problem in MTL is particularly challenging due to the exponential number of possible task combinations (Aribandi et al., 2021; Fifty et al., 2021; Standley et al., 2020; Song et al., 2022). What makes it worse for the exhaustive search is that each trial involves a complete training and evaluation procedure, leading to computational and optimization burden. Moreover, inappropriate task groupings may result in performance degradation due to negative transfer between tasks (Sener & Koltun, 2018; Wu et al., 2020; Maninis et al., 2019). Existing solutions have struggled to address these challenges, often suffering from scalability and modularity issues, making their practical application in real-world scenarios nearly infeasible.
@@
In this paper, we propose a novel data-driven method for task grouping in MTL for classification tasks, which overcomes the scalability and modularity limitations. Our method utilizes the concept of Data Maps (Swayamdipta et al., 2020), hand-crafted features that capture the training behavior of each classification task during MTL training. By analyzing these data maps, we can identify task groupings, both hard and soft ones, that promote positive transfer and mitigate negative interference as much as possible. We demonstrate the effectiveness of our method through extensive experimentation, including experiments on an unprecedented number of tasks, scaling up to 100 tasks to emphasize the practicality of our approach.
The contributions of this paper can be summarized as follows:
@@
We propose a novel data-driven method for task grouping in MTL, addressing the challenges of scalability and modularity.
We propose a mechanism that utilizes our soft-grouping results, enabling model specialization via loss weighting.
@@ 2 Related Work
MTL has been extensively studied to leverage the benefits of information sharing among related tasks, which can serve as an inductive bias to improve modeling performance (Caruana, 1997; Zhang & Yang, 2022). Another perspective on MTL is that it enables more efficient utilization of the model capacity by focusing on learning relevant features and reducing the impact of irrelevant signals, which contributes to overfitting, leading to better generalization. However, when tasks lack shared information, they compete for the limited model capacity, resulting in performance degradation (Sener & Koltun, 2018; Wu et al., 2020; Maninis et al., 2019). To address this challenge, task grouping has emerged as a promising solution to identify subsets of tasks that can be trained together, avoiding negative interference and promoting improved performance.
@@
Traditionally, the decision of task grouping has been approached through costly cross-validation techniques or human expert knowledge (Zhang & Yang, 2022). However, these methods have limitations when applied to different problem domains and may not scale well. Some attempts have been made to approach the problem differently enabling the models to automate the search over which parameters to share among particular tasks (Zhang et al., 2022a; Misra et al., 2016). Methods such as Neural Architecture Search (Liu et al., 2018; Huang et al., 2018; Chen et al., 2023; Zhang et al., 2022a; Sun et al., 2020; Vandenhende et al., 2019), Soft-Parameter Sharing (Ruder et al., 2019; Long et al., 2017; Misra et al., 2016), and asymmetric information transfer (Lee et al., 2016, 2018; Huang et al., 2022a) have been developed. However, these models often exhibit poor generalization and struggle to perform well on diverse tasks and domains. Besides, they often require a large model capacity and do not thus scale well with a large number of tasks.
@@
Therefore, gradient-based approaches (Fifty et al., 2021; Strezoski et al., 2019) have also been explored to determine task grouping in advance. The Task Affinity Grouping (TAG) approach (Fifty et al., 2021), which leverages gradients to determine task similarity, is an example of such an approach. Nevertheless, it has complex training paradigm and requires more forward and backward passes to compute the inter task affinities, putting an issue with scalability even if we enhance the solution’s modularity. Another method, called Higher-Order Approximation (HOA) (Standley et al., 2020), reduces the exponential number of MTL training, from the exhaustive search, by considering only the quadratic pairs of task combinations. However, even with such relaxation, the scalability of HOA remains limited, particularly when dealing with a large number of tasks.
@@
The task grouping problem has been addressed in recent studies through a Meta-Learning approach (Song et al., 2022), aiming to create a meta-learner that can estimate task grouping gains. Nevertheless, the computational demands of this approach pose practical challenges for real-world applications; it requires training MTL networks for every chosen task combination in the training set for multiple iterations. It furthermore outputs all the possible gains of every task combination, whose numbers grow exponentially, and runs a search algorithm over these exponentially growing gains to find the optimal grouping. As a result, the scalability of this solution is severely limited, making it less feasible for a larger number of tasks.
@@ 3.1 Preliminaries
Notations In our paper, we use the following notations consistently. The set of all tasks is denoted as , where represents the number of tasks and . The total number of training data points is denoted as . We calculate the data maps at specific epochs, and the set of epochs is represented as , where corresponds to the epoch. The task clusters are denoted by , and each cluster has an associated centroid . The participation of each task in cluster is represented by , with the constraint that , indicating the percentage of membership; is the weight vector of the tasks in cluster . The values of range from to , where signifies full membership and indicates no membership.
MTL Architecture The MTL procedure for a given task combination, consisting of tasks denoted as , is defined as training with a joint objective for these tasks (Equation 1) where is the accumulated loss value of the cluster, is the loss of the task, and is an optional task weight of the task.
Following the previous approaches (Fifty et al., 2021; Standley et al., 2020; Song et al., 2022), we utilize a commonly employed hard-sharing multi-head architecture (Figure 1) for all our MTL experiments, where a single feature extractor is used to obtain shared representations, and separate task-specific heads are employed to output the result. Additionally, for all the experiments, we maintain the same data splits, via prior seeding, and keep the optimization algorithm and other hyperparameters fixed; this is to make sure any variability in the performance is only attributed to the task grouping and the corresponding weights if any.
@@ 3.2 Data Maps as Task Features
Data Maps (Swayamdipta et al., 2020), originally developed as a model-based tool for characterizing and diagnosing NLP datasets, serve as a valuable component in our approach. They leverage the model behavior concerning individual data point instances of the training set for each task. In our work, we employ Data Maps as task features due to their simplicity, scalability, and ability to extract them on the fly without prior knowledge of the model architecture, thus enhancing the modularity of our approach.
The concept behind Data Maps revolves briefly around extracting two essential values for each data point: the model confidence () of the true class, which is the average probability of the true class over some epochs, and the variability () of this confidence, which is the standard deviation of the true class probabilities over the same epochs. For a particular task, the data map shape is where is the training size. Figure 2 shows an example of the resulting Data Map for an example task extracted from CIFAR10 dataset (Krizhevsky et al., 2009).
@@
In the extraction of data maps, we employ two approaches. The first approach involves building a single MTL model that incorporates all tasks and extracting the data maps directly from this unified model. Alternatively, we utilize the second approach, where individual models are constructed for each task, resulting in multiple STLs, and merging the data maps obtained from each model. Our results are primarily based on the first approach, as it offers the advantage of a single training procedure, simplifying computational complexity, and streamlining experimentation, while having the same qualitative results as the STL.
@@ 3.3 Task Clustering
With the extracted data maps in hand, our next step is to group the tasks into clusters based on their similarity. We propose three distinct approaches for task clustering: soft clustering, hard clustering, and point-based soft clustering.
In both hard and soft clustering, we represent each task as a vector by concatenating the corresponding data maps. In the case of hard clustering, we employ the k-means algorithm (Lloyd, 1982) to cluster these task vectors, aiming to identify distinct clusters of tasks. To introduce a more nuanced representation of task similarities, we incorporate a modified version of the fuzzification step (Equation 3), from (Bezdek et al., 1984), into our approach, which enables soft clustering, where represents the task vector of the corresponding data maps and represents the fuzzification index. This fuzzification process assigns soft memberships to tasks, allowing for more flexible and comprehensive clustering results. We predominantly rely on the soft clustering approach due to its effectiveness and reliability.
In early experiments, though, we also explored a point-based clustering approach to determine the participation membership of tasks. This approach involves clustering each instance point per task within each data map. Each data point then serves as a vote for its corresponding task, and the participation membership of each task is calculated based on the percentage of data points within the cluster (Equation 4) where represents the data point of the data map taken at epoch for task .
However, we do not heavily rely on this point-based approach in our method. This is because it treats the data points within a data map in isolation, failing to capture the abstract behavior specific to each task and overlooking the evolution of data maps over different epochs.
@@ 3.4 Model Specialization through Loss Weighting
In order to assess the effectiveness of our task grouping results, we use loss weighting as a method of model specialization. We construct MTL models that are tailored to specialize in specific sets of tasks based on the membership weights obtained from soft clustering results. For each cluster, we build an individual MTL model that focuses on the tasks assigned to that cluster according to their corresponding weights (Equation 1). To evaluate the performance of our solution, we apply the weighted average of the models’ outputs according to the membership weights as in Equation 5, where is our final output, is the output according to the cluster, and is the weight vector of the cluster. Figure 3 provides an overview of these operations while inferring the output. By comparing the resulting values with both STL and traditional MTL schemes, we can gain insights into the benefits and improvements brought by our task grouping approach.
@@ 3.5 Theoretical Scalability Comparison
In the theoretical scalability comparison, we evaluate our method against existing literature, focusing on the number of models required for clustering. Table 1 presents the comparison, where lower numbers indicate better scalability. Our approach stands out with excellent scalability, as it only necessitates training a single MTL model to extract data maps and perform clustering, or if we consider extracting data maps from STL models. This offers the most promising scalability potential for a larger number of tasks. That is why we can scale our experiments to a very large number of tasks as in Section 4.
Notice TAG requires one single MTL training, yet this is a customized training procedure where each epoch is effectively processed , in particular, times to compute the inter-task affinities pairs, which like the other methods limits its scalability. Therefore, one MTL training of TAG utilizes the same compute of MTL models trained within the other methods normally.
Furthermore, our method’s data map computation is performed on the fly, making it both model and task agnostic. This feature enhances the modularity of our approach, enabling effortless adaptation to different model architectures and tasks without manual intervention.
| Method | Effective MTL models |
@@ 4 Experiments
In this section, we present a comprehensive overview of our experiments, focusing on assessing the effectiveness of our method and presenting the corresponding results. Section 4.1 outlines the specifics of the datasets utilized in our experimentation, as well as the tasks employed. In Section 4.2, we delve into the details of the model architecture and the hyperparameters used during experimentation. The outcomes of the soft clustering of tasks are presented in Section 4.3, where we highlight the effectiveness of our approach in grouping related tasks. Finally, in Section 4.4, we evaluate the quality of the obtained clustering results comparing them to STL and MTL results.

### v2

@@ Abstract
Multi-Task Learning (MTL) is a powerful technique that has gained popularity due to its performance improvement over traditional Single-Task Learning (STL). However, MTL is often challenging because there is an exponential number of possible task groupings, which can make it difficult to choose the best one because some groupings might produce performance degradation due to negative interference between tasks. That is why existing solutions are severely suffering from scalability issues, limiting any practical application. In our paper, we propose a new data-driven method that addresses these challenges and provides a scalable and modular solution for classification task grouping based on a re-proposed data-driven features, Data Maps, which capture the training dynamics for each classification task during the MTL training. Through a theoretical comparison with other techniques, we manage to show that our approach has the superior scalability. Our experiments show a better performance and verify the method’s effectiveness, even on an unprecedented number of tasks (up to 100 tasks on CIFAR100). Being the first to work on such number of tasks, our comparisons on the resulting grouping shows similar grouping to the mentioned in the dataset, CIFAR100. Finally, we provide a modular implementation111https://github.com/ammarSherif/STG-MTL for easier integration and testing, with examples from multiple datasets and tasks.
@@ 1 Introduction
Multi-Task Learning (MTL) has emerged as a powerful technique in deep learning [39, 7] that allows for joint training of multiple related tasks, leading to improved model performance compared to traditional Single-Task Learning (STL). By leveraging shared representations and knowledge across tasks, MTL enhances generalization and mitigates overfitting. Furthermore, MTL promotes faster learning of related tasks and alleviates the computational requirements of deep learning, making it particularly valuable in scenarios with limited task-specific data. That is why MTL has gained significant attention in various domains, including computer vision [8, 25, 30], natural language processing [41, 26, 16, 4], speech recognition [13, 40], and healthcare [26, 2, 15], and has shown promising results in improving accuracy, robustness, and efficiency. However, effectively harnessing the potential of MTL poses several challenges, including the identification of optimal task groupings [29, 9, 30] and the management of negative interference between tasks [28, 35, 24].
@@
The task grouping problem in MTL is particularly challenging due to the exponential number of possible task combinations [1, 9, 30, 29]. What makes it worse for the exhaustive search is that each trial involves a complete training and evaluation procedure, leading to computational and optimization burden. Moreover, inappropriate task groupings may result in performance degradation due to negative transfer between tasks [28, 35, 24]. Existing solutions, which address these challenges, often suffer from scalability and modularity issues, making their practical application in real-world scenarios nearly infeasible.
@@
In this paper, we propose a novel data-driven method for task grouping in MTL for classification tasks, which overcomes the scalability and modularity limitations. Our method utilizes Data Maps [33], data-driven features that capture the training dynamics of each classification task during an MTL training. By analyzing these data maps, we can identify task groupings, both hard and soft ones, that promote positive transfer and mitigate negative interference as much as possible. We demonstrate the effectiveness of our method through extensive experimentation, including experiments on an unprecedented number of tasks, scaling up to 100 tasks to emphasize the practicality of our approach, where the generated groupings aligned with the dataset categorization. We also illustrate the performance improvement of the method in comparison to using both MTL and STL for training. Furthermore, we emphasize the practicality of our approach by providing a modular code implementation, making it easier for researchers and developers to integrate and test our method on their own datasets and tasks.
The contributions of this paper are summarized below:
@@
We propose a novel data-driven method for classification task grouping in MTL, addressing the challenges of scalability and modularity, by re-proposing the usage of data maps as task features.
We utilize soft-clustering weights to enable model specialization via loss weighting.
@@
We provide a modular code implementation of our method, facilitating its adoption and usage by both the research and the industry communities.
@@ 2 Related Work
MTL has been extensively studied to leverage the benefits of information sharing among related tasks, which can serve as an inductive bias to improve modeling performance [5, 39]. Another perspective on MTL is that it enables more efficient utilization of the model capacity by focusing on learning relevant features and reducing the impact of irrelevant signals, which contributes to overfitting, leading to better generalization. However, when tasks lack shared information, they compete for the limited model capacity, resulting in performance degradation [28, 35, 24]. To address this challenge, task grouping has emerged as a promising solution to identify subsets of tasks that can be trained together. This helps with avoiding negative interference and promoting improved performance.
Traditionally, the decision of task grouping has been approached through costly cross-validation techniques or human expert knowledge [39]. However, these methods have limitations when applied to different problem domains and do not scale well. Therefore, some simply used correlation, either between model task predictions [36], data labels, or task weights [37], to recognize related tasks. For instance, in medical imaging, these methods have been applied to improve predictions for Cardiac Indices [36, 37]. However, these approaches are ineffective in scenarios lacking explicit correlations between task labels, in the dataset itself, such as exclusive membership binary classification tasks. Furthermore, both works aim to know task relationships to improve the model predictions of Cardiac Indices in multi-view or multi-modal settings, using a single model. This is quite different from our objective of finding the relations to split the tasks into different models to avoid negative interference.
@@
Other attempts have been made to approach the problem differently enabling the models to automate the search over which parameters to share among particular tasks [38, 25]. Methods such as Neural Architecture Search [20, 12, 6, 38, 32, 34], Soft-Parameter Sharing [27, 23, 25], and asymmetric information transfer [18, 19, 11] have been developed. However, these models often exhibit poor generalization and struggle to perform well on diverse tasks and domains. Besides, they often require a large model capacity and do not thus scale well enough with a large number of tasks.
@@
Therefore, gradient-based approaches [9, 31, 14] have also been explored to determine task grouping in advance. The Task Affinity Grouping (TAG) approach [9], which leverages gradients to determine task similarity, is an example of such an approach. Nevertheless, it has complex training paradigm and requires more forward and backward passes to compute the inter task affinities, putting an issue with scalability even if we enhance the solution’s modularity. Another method, called Higher-Order Approximation (HOA) [30], reduces the exponential number of MTL training, from the exhaustive search, by considering only the quadratic pairs of task combinations. However, even with such relaxation, the scalability of HOA remains limited, particularly when dealing with a large number of tasks.
Recent studies have embraced the utilization of accumulated gradients to depict individual tasks as vectors [14], offering improved scalability by necessitating fine-tuning of only models. However, this method encounters significant limitations concerning modularity. The requirement to fine-tune a pre-trained model for each task results in a dependency where all tasks must share the same pre-trained model, which could pose challenges when dealing with a large number of tasks. It has also a constraint that all tasks must share the same architecture because task vectors are expressed in the model’s weight space. Furthermore, the work focuses on using task vectors to enhance fine-tuning and transfer learning, rather than grouping the tasks into different MTL models to mitigate negative interference.
@@
The task grouping problem has as well been addressed differently through a Meta-Learning approach [29], aiming to create a meta-learner that can estimate task grouping gains. Nevertheless, the computational demands of this approach pose practical challenges for real-world applications; it requires training MTL networks for every chosen task combination in the training set for multiple iterations. It furthermore outputs all the possible gains of every task combination, whose numbers grow exponentially, and runs a search algorithm over these exponentially growing gains to find the optimal grouping. As a result, the scalability of this solution is severely limited, making it less feasible for a larger number of tasks.
@@ 3.1 Preliminaries
Notations In our paper, we use the following notations. The set of all tasks is denoted as , where represents the number of tasks and . The total number of training data points is denoted as . We calculate the data maps at specific epochs, and the set of epochs is represented as , where corresponds to the epoch and . The task clusters are denoted by , and each cluster has an associated centroid and . The association weight, soft-cluster weight, of each task to a cluster is represented by , with the constraint that ; are the weight vector and weighted loss of all the tasks in cluster respectively, while simply represents the resulting loss of model denoting cluster . Notice that the values of range from to , where signifies full membership and indicates no membership.
MTL Architecture Following the previous approaches [9, 30, 29], we utilize a commonly employed hard-sharing multi-head architecture (Figure 1) for all our MTL experiments, where a single feature extractor is used to obtain shared representations, and separate task-specific heads are employed to output the result. Additionally, for all the experiments, we maintain the same data splits, via prior seeding, and keep the optimization algorithm and other hyperparameters fixed; this is to make sure any variability in the performance is only attributed to the task grouping and the corresponding weights if any.
@@ 3.2 Data Maps as Task Features
Data Maps [33] has been originally developed as a model-based tool for characterizing and diagnosing NLP datasets. We re-propose using them as task features during training in an MTL settings given their ability to capture the model behavior of each individual training data point with respect to each task, which is a valuable characterization to the task itself. For that reason, we employ them, in our work, as task features due to their simplicity, scalability, and ability to extract them on the fly without prior knowledge of the model architecture, thus enhancing the modularity of our approach.
The concept behind Data Maps revolves briefly around extracting two values for each data point: the model confidence () of the true class, which is the average probability of the true class over the epochs, and the variability () of this confidence, which is the standard deviation of the true class probabilities over the same epochs. For a particular task, the data map shape is . Figure 2 shows a visualization of the resulting Data Map for an example task extracted from CIFAR10 dataset [17].
@@
In the extraction of data maps, we could have used two approaches. The first approach involves building a single MTL model that incorporates all tasks and extracting the data maps directly from this unified model. Alternatively, we can utilize the second approach, where individual models are constructed for each task, resulting in multiple STLs, and merging the data maps obtained from each model. Our results are primarily based on the first approach, as it offers the advantage of using a single training model, simplifying computational complexity, enhancing scalability, and streamlining experimentation, while having a comparable results to STL, as shown in Section 4.4.
@@ 3.3 Task Clustering
With the extracted data maps in hand, our next step is to group the tasks into clusters based on their similarity. We propose using soft clustering to obtain class membership weights, . To obtain them, we represent each task as a vector by flattening the concatenated data maps, extracted at different epochs. We then employ the k-means algorithm [22] to cluster these task vectors, aiming to identify distinct clusters of tasks, into hard clusters. To introduce a more nuanced representation of task similarities, we incorporate the fuzzification step [3] into our approach, yet we modify it for better computational efficiency, equation 2. This enables soft clustering, where represents the task vector of the corresponding data maps and represents the fuzzification index. This fuzzification process assigns soft memberships to tasks, allowing for more flexible and comprehensive clustering results. We predominantly rely on the soft clustering approach due to its effectiveness and reliability in addition to allowing model specialization via loss weighting, as explained in Section 3.4.
@@ 3.4 Model Specialization through Loss Weighting
In order to assess the effectiveness of our task grouping results, we use loss weighting as a method of model specialization. We construct MTL models that are tailored to specialize in specific sets of tasks based on the membership weights obtained from soft clustering results. For each cluster, we build an individual MTL model that focuses on the tasks assigned to that cluster according to their corresponding weights (Equation 3). To evaluate the performance of our solution, we apply the weighted average of the models’ outputs according to the membership weights as in Equation 4, where is the output of the model and represents Hadamard, element-wise, product. Figure 3 provides an overview of these operations while inferring the output. By comparing the resulting values with both STL and traditional MTL schemes in Section 4.4, we can gain insights into the benefits and improvements brought by our task grouping approach.
Final Output | (4) |
@@ 3.5 Theoretical Scalability Comparison
In the theoretical scalability comparison, we evaluate our method against existing literature, focusing on the required number of trained models to get the clustering results. Table 1 presents the comparison, where lower numbers indicate better scalability. Our approach stands out with excellent scalability, as it only necessitates training a single MTL model to extract data maps and perform clustering, or even if we consider extracting data maps from STL models. This offers the most promising scalability potential for a larger number of tasks. That is why we can scale our experiments to a very large number of tasks as in Section 4. We also highlight the note of Table 1 that although TAG requires a single model to extract the groupings, we compute the gradients , in particular, times than usual to compute the inter-task affinities pairs. Therefore, one model training of TAG utilizes the same asymptotic time of normally trained models.
Furthermore, our method’s data map computation is performed on the fly, making it both model and task agnostic. This feature enhances the modularity of our approach, enabling effortless adaptation to different model architectures and tasks without manual intervention. Refer to Appendix B for further discussion on the modularity aspect of our method.
Method | Number of Models |
@@
TAG originally requires one model, yet we do updates to compute the inter-task affinities. For that reason a single training model of TAG is asymptotically equivalent to training models from other approaches.
@@ 4 Experiments
In this section, we present a comprehensive overview of our experiments, focusing on assessing the effectiveness of our method and presenting the corresponding results. As outlined in Section 3.5, the methods in literature has severe scalability issues. This makes a direct comparison of performance across all methods unfeasible, considering their poor scalability, the substantial number of tasks involved in our evaluation, and our limited computational power. To illustrate, the clustering results of just tasks for HOA necessitate the training of models. Therefore, we evaluate our method against both the MTL and STL results to showcase the added information gained from our clustering results.
Section 4.1 outlines the specifics of the datasets utilized in our experimentation, as well as the tasks employed. In Section 4.2, we delve into the details of the model architecture and the hyperparameters used during experimentation. The outcomes of the soft clustering of tasks are presented in Section 4.3, where we highlight the effectiveness of our approach in grouping related tasks. Finally, in Section 4.4, we evaluate the quality of the obtained clustering results comparing them to STL and MTL results. Appendix B illustrates our modular pipeline structure for all the experiments for easier reproducibility.